There are four activities on the critical path. Coincidentally, their standard deviations are all equal to 4. The standard deviation of the critical path is therefore equal to:

  • 2.

  • 4.

  • 8.

  • 16.

  • 64.
Answer #1

- Variance = (Standard deviation)2 Variance a critical path = Vane + Vos 2 + Vars + Valu =42+42 +4² +42 = 64 Standard deviati